OpenStudy (foreverandeverandeverrrr021014):

Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). please help answers are: A.-7 B.-3 C.10 D.33 would it be -3?

OpenStudy (studygurl14):

No. Try plugging in 2 for x in g(x). Take what you get for that, and then plug it in for x in f(x)

OpenStudy (dinamix):

thr right answer is A , i 'm sure 200%

OpenStudy (dinamix):

@foreverandeverandeverrrr021014 did u see

OpenStudy (foreverandeverandeverrrr021014):

imn not quite sure how you got 7

OpenStudy (dinamix):

g(2)=-17 f(g(2) = -17+8 =-7

OpenStudy (foreverandeverandeverrrr021014):

Let f(x) = −4x + 7 and g(x) = 10x − 6. Find f(g(x)). a.−40x + 24 b. −40x + 31 c.−40x + 64 d. −40x + 70 what about thisone?

OpenStudy (dinamix):

f(g(x))= -4(10x-6)+7 = -40x+24+7 =-40x+31 @foreverandeverandeverrrr021014

OpenStudy (foreverandeverandeverrrr021014):

oh thanks! your more help than the lesson on this

OpenStudy (dinamix):

np mate dude u are welcome